# Optimizing Temperature for Language Models with Multi-Sample Inference

### v1

@@ Abstract
Multi-sample aggregation strategies, such as majority voting and best-of-N sampling, are widely used in contemporary large language models (LLMs) to enhance predictive accuracy across various tasks. A key challenge in this process is temperature selection, which significantly impacts model performance. Existing approaches either rely on a fixed default temperature or require labeled validation data for tuning, which are often scarce and difficult to obtain. This paper addresses the challenge of automatically identifying the (near)-optimal temperature for different LLMs using multi-sample aggregation strategies, without relying on task-specific validation data. We provide a comprehensive analysis of temperature’s role in performance optimization, considering variations in model architectures, datasets, task types, model sizes, and predictive accuracy. Furthermore, we propose a novel entropy-based metric for automated temperature optimization, which consistently outperforms fixed-temperature baselines. Additionally, we incorporate a stochastic process model to enhance interpretability, offering deeper insights into the relationship between temperature and model performance.
@@
Figure 1: (a) The entropy turning point (EntP) is defined as the point where the log-scale of token-level entropy shifts from concave to convex. (b) The accuracy tested in EntP is highly correlated with the best accuracy from grid search.
@@ 1 Introduction
A common practice in prior evaluations is to use the same temperature across all methods despite variations in training datasets, task types, model sizes, and aggregation strategies. This is clearly suboptimal. An alternative approach is to empirically tune the temperature using labeled validation data for each task, dataset, model size, and aggregation strategy (Zhang et al., 2024a; Dhuliawala et al., 2024). However, such a process is tedious and time-consuming and heavily dependent on the availability of labeled validation data, limiting its applicability when such data are scarce.
@@ 2 Preliminary & Related Work
Before moving to our main contributions, we first review how language models typically generate samples and an introduction to multi-sample aggregation strategies.
@@
Language models typically generate outputs for generative tasks by autoregressively sampling from the conditional probability distribution over the next token, given both the input context and previously generated tokens. Formally, for an input $X$ and an output sequence $Y=(y_{1},y_{2},\dots,y_{N})$, the probability of producing $Y$ is given by:
@@
    where $\mathbb{I}(\cdot)$ is the indicator function, which returns 1 if its argument is true and 0 otherwise. This method is frequently used in where evaluating whether two outputs are equivalent is relatively easy. The method is also called self-consistency (Wang et al., 2022).
@@ 3.2 Correlation Between Training-Task Similarity and Optimal Temperature
Our goal is to establish a general relationship between a model’s learned distribution and its optimal temperature for a task. Our key intuition is that token-level entropy can serve as a surrogate for a model’s ‘distance’ from a target task and that this distance helps identify the optimal temperature.
@@
In practice, this suggests using a higher temperature (e.g., $T=0.9\sim 1.1$) for task-finetuned models and a lower temperature (e.g., $T=0.5\sim 0.7$) for more general-purpose models (pretrained or instruction-finetuned).
Figure 4: Entropy Curve Characteristics. (a) The token-level entropy $\mathcal{H}$ (solid blue line) increases slowly at lower temperatures and then jumps sharply at a critical turning point. In contrast, the entropy for a fixed (greedy) generation stays low (dotted blue line). $\log(\mathcal{H})$ (red line) reveals a transition from concavity to convexity that aligns with the sharp increase in $\mathcal{H}$, marking the entropy turning point (EntP). (b) EntP hits the best temperature, and it varies between different models.
@@ 4.2 Turning Point Temperature Selection (TURN)
Given the token-level entropy curve of a language model on a specific task, how can we identify a suitable temperature for multi-sample aggregation strategies? Inspired by the difference in the shapes of the entropy curve: When the temperature remains low, the entropy increases flatly. However, when the sampling temperature is near the spike, the entropy increases (super)-exponentially, implying a quality drop in samples. Therefore, after taking the logarithm of the entropy curve (shown in Figure 4(a), red line), the flat part becomes concave while the exponentially-increase part becomes convex. We define the entropy turning point (EntP) as the temperature where the log entropy curve becomes convex. Figure 4(b) tests the llemma-7b base model and its task-finetuned variant [^2] Model link: https://huggingface.co/ScalableMath/llemma-7b-sft-metamath-level-1to3-hf (Sun et al., 2024), and EntP matches the position with the highest accuracy and varies between different models. Based on EntP, we develop a new method for automatic temperature prediction in multi-sample aggregation strategies, called Turning Point Temperature Selection (TURN).
The optimal temperature should be around EntP to achieve both sample quality and diversity. At the same time, we found that some aggregation methods may be more tolerant of quality drops (e.g., for best-of-N, only one sample is enough to be correct). So we added a small adaptation factor $\beta$ based on the aggregation function, and it is set to 00 and $+0.1$ for majority voting and best-of-N, respectively. The aggregation adaptation for best-of-N is calculated in the MATH dataset but can be directly applied to other tasks. Refer to Appendix C for details.
Specifically, given a language model $\mathcal{M}$, a task $\mathcal{T}=\{X_{1},\ldots,X_{k}\}$ with $k$ input instances, and an aggregation method $\mathcal{A}$. To estimate the token-level entropy, we random sample $N$ times. In each time, we randomly choose an input instance $X_{i}$, and generate one sample by $\mathcal{M}$ under each candidate temperature $t_{j}=j\cdot t$ (with $t$ being the temperature interval and $j=0,1,\ldots,J$, where $J=\lfloor t_{\max}/t\rfloor$). These entropies are then aggregated to calculate the average entropy $\mathcal{H}(t_{j})$ at each temperature $t_{j}$. By taking the logarithm, we obtain $\ell(t_{j})=\log\mathcal{H}(t_{j})$.
Next, we identify EntP index $j^{*}$, where the second derivative of $\ell$ changes from negative to positive and select its corresponding temperature $j^{*}\cdot t$. Then we add the aggregation adaptation factor $\beta$ to form the final prediction. The pseudocode for our algorithm is listed in Algo. 1.
@@
Figure 5: Stochastic Process Model. We run our process model in the setting: $N_{0}=10$, $N_{1}=30000$, $L_{0}=0$, $\sigma_{0}=1$, $L_{1}=-10$, and $\alpha=2$. (a) The entropy curve is similar to that of the real language model: flat at first, and then sharply increases. (b) We calculate the relation between temperature and the percentage of improper tokens in simulation, and the percentage of improper tokens quickly increases after EntP.
@@ 5 Evaluating TURN
Table 1: The prediction from our algorithm TURN, the optimal temperature ranges from grid search, and the performance drop (PD) for various models tested in the MATH and MBPP datasets. TURN achieved hit rates of $12/13$ and $11/13$, average temperature gaps of $0.023$ and $0.015$, and average performance drop of $0.32\%$ and $0.59\%$.
MATH Dataset with Majority Voting
Model Predict Optimal Range PD($\downarrow$)
@@
MBPP Dataset with Best-of-N
Model Predict Optimal Range PD($\downarrow$)
@@
Qwen2.5-Coder-7B-Instruct 0.9 0.7–1.10 0%
@@
MATH Sum(Gap) ($\downarrow$) 4.6 2.0 0.4 0.4 2.0 3.6 0.3 0.3
Mean(Drop) ($\downarrow$) 8.6% 3.5% 1.0% 0.8% 3.1% 7.2% 0.3% 0.3%
MBPP Sum(Gap) ($\downarrow$) 8.2 5.6 3.0 0.8 0.2 1.2 0.2 0.6
Mean(Drop) ($\downarrow$) 22.5% 10.7% 5.1% 1.5% 0.9% 4.2% 0.5% 1.1%
Average Sum(Gap) ($\downarrow$) 6.4 3.8 1.7 0.6 1.1 2.4 0.25 0.45
Mean(Drop) ($\downarrow$) 15.55% 7.1% 3.05% 1.15% 2.0% 5.7% 0.4% 0.7%
@@ 5.1 Experiment Setup
Table 3: Variance of Entropy Estimation. We report the average variance of the entropy curve and the variance of estimated temperature $T_{\text{pred}}$ under different sample sizes with $50$ trails on Llama3.1-8B-Instruct on MATH. A small sample size (e.g., $40$) is sufficient for entropy estimation in TURN for its low prediction variance and small performance drop.
@@ 5.2 Evaluation Metrics
  • Performance Drop (PD): The accuracy loss compared to the best temperature found via grid search.
@@ 5.3 Baseline
As no existing method automatically adjusts temperatures in multi-sample aggregation strategies, we compare against a fixed temperature baseline. We search over $\{0.1,0.3,0.5,0.7,0.9,1.1\}$ and select the temperature that maximizes overall accuracy. This mimics a common yet suboptimal practice where developers apply a single temperature across all models, disregarding variations in model behavior and task requirements.
@@ 5.4 Results
Number of Samples for Temperature Estimation: Finally, we assess the efficiency of TURN by examining the prediction variance under different sample sizes for Llama-3.1-8B-Instruct on MATH. As shown in Table 3, we report the average variance of the entropy curve across all choices of $T$, the variance of predicted temperature, and the average performance drop. We find out that even with a moderate sample size (e.g. $40$ samples), the variance remains low and the performance drop is tiny (0.2%), suggesting that a small sampling budget is sufficient for accurate temperature estimation and thus proves the efficiency of our algorithm.
@@ A.3 Metric Calculation
To compute majority-vote results for the MATH dataset, we consider two samples to have the same answer if they match after normalization. For the pass@K metric, we follow the definition in Chen et al. (2021). Let $N$ be the total number of samples and $C$ be the number of correct samples. Then $\mathrm{pass}@K$ is defined as:
@@ B.1.2 Interpreting the Error Rate
Define a random variable $E_{t}$ that indicates whether an error occurred at time $t$:
@@ B.1.3 Error Rate Update Rules
After each step, the error rate $x_{t+1}$ is updated based on whether the token at time $t$ was improper or proper:
@@
The error rate is increased. Intuitively, making a mistake can make the model more likely to continue making errors. Formally, we update:
@@
For $x_{t}\in[0,1]$ and $\alpha>1$ ($\alpha$ is a hyperparameter), it can be seen that $x_{t+1}\geq x_{t}$. Here $\alpha$ can be regarded as the noise tolerance rate, measuring how stable the model is when suffering from unexpected noises, and we will try different $\alpha$ in experiments.
@@
The error rate is reduced, reflecting a ‘reinforcement’ of correct behavior. We do this by:
@@
It generally makes it smaller, so this step lowers the error rate. However, we do not allow the error rate to drop below the initial baseline $x_{\text{init}}$.
@@ B.1.4 Linking Initial Error Rate and Temperature
At step $t$, the model’s token probability mass is divided into:
  • Improper tokens with total probability $P_{1,\text{improper}}=x_{t}$.
@@
Therefore, by definition:
$x_{\text{init}}=P_{1,\text{improper}}.$
Under higher temperatures $T$, the softmax distribution flattens, increasing $P_{1,\text{improper}}$ because the number of improper tokens is large but their logits are low. Thus, $x_{\text{init}}$ increases as $T$ increases.
@@ B.1.5 Type of Tokens during decoding
  • Proper tokens: A small number of tokens with high logits. Let $N_{0}$ be the number of proper tokens. To capture the multi-peak behavior, the logits of proper tokens $l_{0,1},...,l_{0,N_{0}}$are sampled from the Gaussian distribution $\mathcal{N}(L_{0},\sigma_{0})$.
  • Low Probability Improper tokens: Many low-logit tokens where language models seldom choose them. Let $N_{1}$ be the number of tokens in this type and their logits are set to $L_{1}$ for simplicity.
  • High Probability Improper tokens: Due to insufficient training or errors in training data, some tokens may have exceptionally high logits but are logically improper in (e.g., the token $3$ in $1+1=3$). Since different language models behave differently and this type of token will be selected regardless of temperature, we only consider decoding without high-probability improper tokens in our discussion.
For the first two types of tokens, we have:
$L_{0}>L_{1},\ \ N_{0}\ll N_{1}.$
@@ B.1.6 Token Probability During Sampling
At each step $t$, the probability of producing improper tokens is $x_{t}$. Let $p_{t,\text{proper/improper},j}$ be the probability of the $j$-th proper/improper tokens at step $t$. For the improper tokens, we have:
@@
This ensures that the relative ordering of probabilities for the proper tokens remains determined by their logits, while the total mass allocated to improper tokens is $x_{t}$.

### v2

@@ Abstract
Multi-sample aggregation strategies, such as majority voting and best-of-N sampling, are widely used in contemporary large language models (LLMs) to enhance predictive accuracy across various tasks. A key challenge in this process is temperature selection, which significantly impacts model performance. Existing approaches either rely on a fixed default temperature or require labeled validation data for tuning, which are often scarce and difficult to obtain. This paper addresses the challenge of automatically identifying the (near)-optimal temperature for different LLMs using multi-sample aggregation strategies, without relying on task-specific validation data. We provide a comprehensive analysis of temperature’s role in performance optimization, considering variations in model architectures, datasets, task types, model sizes, and predictive accuracy. Furthermore, we propose a novel entropy-based metric for automated temperature optimization, which consistently outperforms fixed-temperature baselines. Additionally, we incorporate a stochastic process model to enhance interpretability, offering deeper insights into the relationship between temperature and model performance. Our code is available at https://github.com/StigLidu/TURN.
@@
Figure 1: (a) The entropy turning point (EntP) (green star) is defined as the temperature point where the log-scale of the token-level entropy of generation (red line) shifts from concave to convex, implying the sudden spike in the entropy curve (blue line). (b) The accuracy tested at EntP is highly correlated with the best accuracy from grid search over temperatures on the MATH dataset.
@@ 1 Introduction
A common practice in prior evaluations is to use the same temperature across all methods despite variations in training datasets, task types, model sizes, and aggregation strategies. This practice is clearly suboptimal. An alternative approach is to empirically tune the temperature using labeled validation data for each task, dataset, model size, and aggregation strategy (Zhang et al., 2024a; Dhuliawala et al., 2024). However, such a process is tedious and time-consuming and heavily dependent on the availability of labeled validation data, limiting its applicability when such data are scarce.
@@ 2 Preliminary & Related Work
Before moving to our main contributions, we first review how language models typically generate samples and provide an introduction to multi-sample aggregation strategies.
@@
Language models typically generate output for generative tasks by autoregressively sampling from the conditional probability distribution over the next token, given both the input context and previously generated tokens. Formally, for an input sequence $X$ and an output sequence $Y=(y_{1},y_{2},\dots,y_{N})$, the probability of producing $Y$ is given by the following:
@@
    where $\mathbb{I}(\cdot)$ is the indicator function, which returns 1 if its argument is true and 0 otherwise. This method is frequently used where evaluating whether two outputs are equivalent is relatively easy. The method is also called self-consistency (Wang et al., 2022).
@@ 3.2 Correlation Between Training-Task Similarity and Optimal Temperature
Our goal is to establish a general relationship between a model’s learned distribution and its optimal temperature for a task. Our key intuition is that token-level entropy can serve as a proxy for how distant a model is from a target task and that this distance helps identify the optimal temperature.
@@
In practice, this suggests using a higher temperature (e.g., $T=$ 0.9–1.1) for task-finetuned models and a lower temperature (e.g., $T=$ 0.5–0.7) for more general-purpose models (pretrained or instruction-finetuned).
Figure 4: Entropy Curve Characteristics. (a) The token-level entropy $\mathcal{H}$ (solid blue line) increases slowly at lower temperatures and then jumps sharply at a critical turning point. In contrast, the entropy for a fixed (greedy) generation stays low (dotted blue line). $\log(\mathcal{H})$ (red line) reveals a transition from concavity to convexity that aligns with the sharp increase in $\mathcal{H}$, marking the entropy turning point (EntP). (b) EntP aligns with the best temperature and varies across different models.
@@ 4.2 Turning Point Temperature Selection (TURN)
Given the token-level entropy curve of a language model on a specific task, how can we identify a suitable temperature for multi-sample aggregation strategies? Inspired by the difference in the shapes of the entropy curve: When the temperature remains low, the entropy increases flatly. However, when the sampling temperature is near the spike, the entropy increases (super)-exponentially, implying a quality drop in samples. Therefore, after taking the logarithm of the entropy curve (shown in Figure 4(a), red line), the flat part becomes concave while the exponentially-increasing part becomes convex. We define the entropy turning point (EntP) as the temperature where the log entropy curve becomes convex. Figure 4(b) tests the llemma-7b base model and its task-finetuned variant [^2] Model link: https://huggingface.co/ScalableMath/llemma-7b-sft-metamath-level-1to3-hf (Sun et al., 2024), and EntP matches the position with the highest accuracy and varies between different models. Based on EntP, we develop a new method for automatic temperature prediction in multi-sample aggregation strategies, called Turning Point Temperature Selection (TURN).
The optimal temperature should be around EntP to achieve both sample quality and diversity. At the same time, we found that some aggregation methods may be more tolerant of quality drops (e.g., for best-of-N, only one sample is enough to be correct). So we added a small adaptation factor $\beta$ based on the aggregation function, and it is set to 00 and $+0.1$ for majority voting and best-of-N, respectively. The aggregation adaptation for best-of-N is calculated in the MATH dataset but can be directly applied to other tasks. Refer to Appendix C for more details.
Specifically, given a language model $\mathcal{M}$, a task $\mathcal{T}=\{X_{1},\ldots,X_{k}\}$ with $k$ input instances, and an aggregation method $\mathcal{A}$. To estimate token-level entropy, we randomly sampled $N$ times. In each time, we randomly choose an input instance $X_{i}$ and generate one sample by $\mathcal{M}$ under each candidate temperature $t_{j}=j\cdot t$ (with $t$ being the temperature interval and $j=0,1,\ldots,J$, where $J=\lfloor t_{\max}/t\rfloor$). These entropies are then aggregated to calculate the average entropy $\mathcal{H}(t_{j})$ at each temperature $t_{j}$. Taking the logarithm, we obtain $\ell(t_{j})=\log\mathcal{H}(t_{j})$.
Next, we identify the EntP index $j^{*}$, where the second derivative of $\ell$ changes from negative to positive, and select its corresponding temperature $j^{*}\cdot t$. Then we add the aggregation adaptation factor $\beta$ to form the final prediction. The pseudocode for our algorithm is listed in Algo. 1.
@@
Figure 5: Stochastic Process Model. We run our process model in the setting: $N_{0}=10$, $N_{1}=30000$, $L_{0}=0$, $\sigma_{0}=1$, $L_{1}=-10$, and $\alpha=2$. (a) The entropy curve is similar to that of the real language model: flat at first, and then sharply increases. (b) We calculate the relation between temperature and the percentage of improper tokens in the simulation, and the percentage of improper tokens quickly increases after EntP.
@@ 5 Evaluating TURN
Table 1: The prediction from our algorithm TURN (Pred.), the optimal temperature ranges (Opt. Range) from grid search, and the performance drop (PD) for various models tested in the MATH and MBPP datasets. TURN achieved hit rates of $12/13$ and $11/13$, average temperature gaps of $0.023$ and $0.015$, and average performance drop of $0.32\%$ and $0.59\%$.
MATH with Majority Voting (Hit Rate: 92.3%)
Model Name Pred. Opt. Range PD ($\downarrow$)
@@
MBPP with Best-of-N (Hit Rate: 84.6%)
Model Name Pred. Opt. Range PD ($\downarrow$)
@@
Qwen2.5-Coder-7B-Instruct 0.9 0.7–1.1 0%
@@
MATH Sum Gap ($\downarrow$) 4.6 2.0 0.4 0.4 2.0 3.6 0.3 0.3
Avg. Drop ($\downarrow$) 8.6% 3.5% 1.0% 0.8% 3.1% 7.2% 0.3% 0.3%
MBPP Sum Gap ($\downarrow$) 8.2 5.6 3.0 0.8 0.2 1.2 0.2 0.6
Avg. Drop ($\downarrow$) 22.5% 10.7% 5.1% 1.5% 0.9% 4.2% 0.5% 1.1%
Average Sum Gap ($\downarrow$) 6.4 3.8 1.7 0.6 1.1 2.4 0.25 0.45
Avg. Drop ($\downarrow$) 15.55% 7.1% 3.05% 1.15% 2.0% 5.7% 0.4% 0.7%
@@ 5.1 Experiment Setup
Table 3: Variance of Entropy Estimation. We report the average variance of the entropy curve and the variance of estimated temperature $T_{\text{pred}}$ under different sample sizes with $50$ trials on Llama3.1-8B-Instruct on MATH. A small sample size (e.g., $40$) is sufficient for entropy estimation in TURN for its low prediction variance and small performance drop.
@@ 5.2 Evaluation Metrics
  • Performance Drop (PD): The accuracy loss compared to the best temperature found by grid search.
@@ 5.3 Baseline
As no existing method automatically adjusts temperatures in multi-sample aggregation strategies, we compare against a fixed temperature baseline. We search for $\{0.1,0.3,0.5,0.7,0.9,1.1\}$ and select the temperature that maximizes the overall accuracy. This mimics a common, yet suboptimal practice where developers apply a single temperature across all models, disregarding variations in model behavior and task requirements.
@@ 5.4 Results
Number of Samples for Temperature Estimation: Finally, we assess the efficiency of TURN by examining the prediction variance under different sample sizes for Llama-3.1-8B-Instruct on MATH. As shown in Table 3, we report the average variance of the entropy curve in all choices of $T$, the variance of the predicted temperature, and the average performance drop. We find that even with a moderate sample size (e.g., $40$ samples), the variance remains low and the performance drop is tiny (0.2%), suggesting that a small sampling budget is sufficient for accurate temperature estimation and thus proves the efficiency of our algorithm.
@@ 6 Conclusion
Acknowledgments
SW would like to thank Convergent Research and the Lean Focused Research Organization and the Microsoft Accelerating Foundation Models Research Initiative.
@@ A.3 Metric Calculation
To compute the majority vote results for the MATH dataset, we consider two samples to have the same answer if they match after normalization. For the pass@K metric, we follow the definition in Chen et al. (2021). Let $N$ be the total number of samples and $C$ be the number of correct samples. Then $\mathrm{pass}@K$ is defined as:
@@ B.1.2 Interpreting the Error Rate
Define a random variable $E_{t}$ that indicates whether an error occurred at time step $t$:
@@ B.1.3 Error Rate Update Rules
After each step, the error rate of the next step $x_{t+1}$ is updated based on whether the token in this step $t$ is proper or not:
@@
The error rate of the next step $x_{t+1}$ is increased. Intuitively, making a mistake can make the model more likely to continue making errors. Formally, we update:
@@
It can be seen that $x_{t+1}\geq x_{t}$ for $x_{t}\in[0,1]$ and $\alpha>1$ ($\alpha$ is a hyperparameter). Here, $\alpha$ can be considered as the noise tolerance rate, measuring how stable the model is when it experiences unexpected noise, and we try different $\alpha$ in experiments.
@@
The error rate of the next step $x_{t+1}$ is reduced, reflecting a ‘reinforcement’ of correct behavior. We do this by:
@@
It generally makes $x_{t+1}\leq x_{t}$ smaller, so this update lowers the error rate. In particular, we do not allow the error rate to drop below the initial baseline $x_{\text{init}}$.
@@ B.1.4 Linking Initial Error Rate and Temperature
At time step $t$, the token probability mass of the model is divided into:
  • Improper tokens with total probability $P_{1,\text{improper}}=x_{t}$;
@@
Therefore, by definition, we have $x_{\text{init}}=P_{1,\text{improper}}.$ Under higher temperatures $T$, the softmax distribution flattens, increasing $P_{1,\text{improper}}$ because the number of improper tokens is large but their logits are low. Thus, $x_{\text{init}}$ increases as $T$ increases.
@@ B.1.5 Type of Tokens during decoding
  • Proper tokens: Small number of tokens with high logits. Let $N_{0}$ be the number of proper tokens. To capture the multi-peak behavior, the logits of proper tokens $l_{0,1},...,l_{0,N_{0}}$are sampled from the Gaussian distribution $\mathcal{N}(L_{0},\sigma_{0})$.
  • Low Probability Improper tokens: Many low-logit tokens where language models seldom choose them. Let $N_{1}$ be the number of tokens of this type, and their logits are set to $L_{1}$ for simplicity.
  • High Probability Improper tokens: Due to insufficient training or mistakes in training data, some tokens may have exceptionally high logits but are logically improper in (e.g., the token $3$ in $1+1=3$). Since different language models behave differently, we only consider decoding without high-probability improper tokens in our discussion.
For the first two types of tokens, we have $L_{0}>L_{1},\ N_{0}\ll N_{1}.$
@@ B.1.6 Token Probability During Sampling
At each step $t$, the probability of producing improper tokens is $x_{t}$. Let $p_{t,\text{proper/improper},j}$ be the probability of the $j$-th proper/improper tokens at step $t$. For improper tokens, we have:
@@
This ensures that the relative order of the probabilities for the proper tokens remains determined by their logits, while the total mass allocated to the improper tokens is $x_{t}$.